OpenStudy (smurfy14):

if the directions say to Use logarithms to solve each equation and the problem is 9^x=45 how do i set it up and solve?

OpenStudy (amistre64):

9^x = 45 log(9^x) = log(45) x * log(9) = log(45) x = log(45) / log(9) ; use change of base here.. x = log9(45)

OpenStudy (smurfy14):

so x would equal 1.732?